OpenStudy (anonymous):

What is the value of y in the equation 2(4y − 1) = 6? 0 1 6 8

OpenStudy (anonymous):

Distribute first

OpenStudy (anonymous):

So multiply everything in the parenthesis by 2...

OpenStudy (anonymous):

8y - 2?

OpenStudy (anonymous):

2 * 4y = ? 2 * -1 = ?

OpenStudy (anonymous):

CORRECT

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Now is the 2 subtracted or added?

OpenStudy (anonymous):

added?

OpenStudy (anonymous):

8y=8

OpenStudy (anonymous):

Yes, it is added to the other side of the equation.

OpenStudy (anonymous):

6 + 2 is what?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

so the answer is one

OpenStudy (anonymous):

1

OpenStudy (anonymous):

is 8 multiplied or divided?

OpenStudy (anonymous):

multiplied?

OpenStudy (anonymous):

8y=6+2

OpenStudy (anonymous):

it is being multilied to the y, however, we need to do the opposite....SO we divide the 8 from both sides.... 

OpenStudy (anonymous):

8y/8=8/8

OpenStudy (anonymous):

so this leaves us with y = 8/8 what is 8 divided by 8?

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

lol its 0

OpenStudy (anonymous):

or 1?

OpenStudy (anonymous):

no, when a number is divided by itself....it is always 1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

y = 1 HOWEVER it is time to double check

OpenStudy (anonymous):

thx guys! i fanned u both but im gonna give @DIDITHEDAD the medal bcuz she/you did a better job at explaining! i have more questions im gonna post in a new question ;)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2(4y − 1) = 6? y=1 put the 1 in for y 2(4(1)-1) = 6 4 * 1 is four 2 (4 - 1) = 6 do what is in the paranthesis first 2 (3) = 6 2 * 3 = 6 2 times 3 is what? it is six.. so 6=6 and that means the answer is....Y=1!